# The Uniqueness of LLaMA3-70B with Per-Channel Quantization: An Empirical Study

### v1

@@ 9.2.2 The latter block has less large outliers
Figure 23, 24, 25, 26, 27, and 28 show the Q and Up matrices in the 10th, 40th, and 79th (last) Transformer block in LLaMA3-70B. The magnitude of the weight outliers is significantly smaller than that in the first block, demonstrating more robustness to W8 per-channel quantization.
@@ 9.3 Weights of Other LLM models
Figure 29, 30, 31, 32 show the weights with maximum quantization error under W8 per-channel quantization in four LLMs, that is, LLaMA2-70B, LLaMA3-8B, LLaMA3.1-405B, and Qwen2-72B. It is worth noting that the maximum quantization error occurs all at the same layer, which is the “K” matrix in the first Transformer block. We can see that the weight outliers in the layer with maximum quantization error is less than 1.0, which is far smaller than that of LLaMA3-70B models, which is usually greater than 90. The difference of 2 orders of magnitude causes the quantization error of the LLaMA3-70Bs to be much more than the rest of the models.

### v2

@@ 9.2.2 The latter block has less large outliers
Figure 32, 33, 34, 35, 36, and 37 show the Q and Up matrices in the 10th, 40th, and 79th (last) Transformer block in LLaMA3-70B. The magnitude of the weight outliers is significantly smaller than that in the first block, demonstrating more robustness to W8 per-channel quantization.
@@ 9.3 Weights of Other LLM models
Figure 38, 39, 40, 41 show the weights with maximum quantization error under W8 per-channel quantization in four LLMs, that is, LLaMA2-70B, LLaMA3-8B, LLaMA3.1-405B, and Qwen2-72B. It is worth noting that the maximum quantization error occurs all at the same layer, which is the “K” matrix in the first Transformer block. We can see that the weight outliers in the layer with maximum quantization error is less than 1.0, which is far smaller than that of LLaMA3-70B models, which is usually greater than 90. The difference of 2 orders of magnitude causes the quantization error of the LLaMA3-70Bs to be much more than the rest of the models.